OpenStudy (anonymous):

Find the area of the region bounded by the x-axis and the graph of the function: f(x)= (eˣ)/(1+eˣ)² We are studying improper integrals if that helps

OpenStudy (anonymous):

first thing to note is that this is an even function, so you can integrate \[\int_0^{\infty} \frac{e^x}{(1+e^x)^2}dx\] and double the result

OpenStudy (anonymous):

why would you double it?

OpenStudy (anonymous):

because you are being asked for \[\int_{-\infty}^{\infty}\frac{e^x}{(1+e^x)^2}dx\]

OpenStudy (anonymous):

and so since it is even we can take the integral from 0 to infinity, and then multiply by 2

OpenStudy (anonymous):

okay I guess I don't understand even integrals

OpenStudy (anonymous):

"even function" not "even integral" even function means \[f(x)=f(-x)\]

OpenStudy (anonymous):

and how do you know its from -∞ to ∞?

OpenStudy (anonymous):

okay and it's even because anything squared will always be positive

OpenStudy (anonymous):

so that what you see to the left of the y - axis is what you see to the right. i.e. it is symmetric with respect to the y axis. so if you want the integral from -5 to 5, say, you can compute the integral from 0 to 5 and double

OpenStudy (anonymous):

okay gotcha

OpenStudy (anonymous):

you know you are integrating from - infinity to + infinity because of the way the question is asked. area above the x - axis and below the curve

OpenStudy (anonymous):

you are not given any bounds other than the x - axis below and the function above

OpenStudy (anonymous):

how do you know it's not below the x-axis?

OpenStudy (anonymous):

because this function is positive

OpenStudy (anonymous):

okay. so now I integrate

OpenStudy (anonymous):

before we start, it is not even because of the "square" it is even because if \[f(x)=\frac{e^x}{(1+e^x)^2}\] then \[f(-x)=\frac{e^{-x}}{(1+e^{-x})^2}\] and then if you do some algebra we get this is the same as \[f(x)=\frac{e^x}{(1+e^x)^2}\]

OpenStudy (anonymous):

so first we find an anti-derivative of \[f(x)=\frac{e^x}{(1+e^x)^2}\]

myininaya (myininaya):

are you sure the function is even satellite?

myininaya (myininaya):

oh wait...

myininaya (myininaya):

ok you may proceed

myininaya (myininaya):

its even

OpenStudy (anonymous):

by a u substitution. you let \[u=1+e^x\]and \[du=e^xdx\] and so you have \[\int\frac{1}{u^2}du=-\frac{1}{u}=-\frac{1}{(1+e^x)^2}\]

OpenStudy (anonymous):

and yes i am sure it is even because \[\frac{e^{-x}}{(1+e^{-x})^2}=\frac{\frac{1}{e^x}}{1+\frac{1}{2e^x}+\frac{1}{e^{2x}}}\] which, if you multiply top and bottom by \[e^{2x}\] gives \[\frac{e^x}{(1+e^x)^2}\]

OpenStudy (anonymous):

wait... theintegral of 1/u².. isn't that 1/2 u^-1?

myininaya (myininaya):

yes i know satellite i already did it

OpenStudy (anonymous):

never mind!

OpenStudy (anonymous):

@shortiesport, no it is just regular old \[-\frac{1}{u}\]

OpenStudy (anonymous):

ok so now we have the anti derivative. now we get to evaluate but remember that \[\int_0^{\infty} \frac{e^x}{(1+e^x)^2}dx\] means \[\lim_{r\rightarrow \infty}\int_0^r \frac{e^x}{(1+e^x)^2}dx\]

OpenStudy (anonymous):

do I always have to find if it's even or odd? i'm confuseed too because this chapter was kinda about the limit as b approaches inifinity

OpenStudy (anonymous):

so we take \[-\frac{1}{(1+e^x)}\] and replace x by r and 0

myininaya (myininaya):

no you don't have to do the even odd thing

OpenStudy (anonymous):

no, sometimes it is even, sometimes it is odd, and most time neither. now comes the integral part, where we evaluate and then take the limit as r goes to infinity

OpenStudy (anonymous):

so do I write this as the limit as b -> ∞ of -1/(1+eˣ)?

OpenStudy (anonymous):

at r you get \[-\frac{1}{1+e^r}\] and at 0 you get \[-\frac{1}{1+e^0}=-\frac{1}{2}\] so \[\lim_{r\rightarrow \infty}\int_0^r \frac{e^x}{(1+e^x)^2}dx=\lim_{r\rightarrow \infty}-\frac{1}{1-e^r}+\frac{1}{2}\]\]

OpenStudy (anonymous):

which pretty clearly is \[\frac{1}{2}\]

OpenStudy (anonymous):

since \[\lim_{r\rightarrow \infty}-\frac{1}{1-e^r}=0\]

OpenStudy (anonymous):

and therefore your "final answer" is 1

OpenStudy (anonymous):

please notice that although you did not need to use the fact that this was even, it was a huge help. you have an integral that goes from - infinity to infinity, so you are going to have to break it up somewhere.

OpenStudy (anonymous):

in other words it is an improper integral two times, you have to compute the limit at both ends, so being even cut the work in half

OpenStudy (anonymous):

okay.. i think i'm going to have to go back through the steps in slow motion lol. I understand why we found the integral of the original problem = -1/(1+eˣ)². I just don't understand what we did after that

OpenStudy (anonymous):

here is a picture of your function. you can see that it is even, that it sits above the x - axis, and that it approaches 0 nice and fast http://www.wolframalpha.com/input/?i=y%3De^x%2F%281%2Be^x%29^2

myininaya (myininaya):

good job satellite you can pass cal 2 again

OpenStudy (anonymous):

like what is the limit even telling me

OpenStudy (anonymous):

ok after that we computed this: \[\int_0^r\frac{e^x}{(1+e^x)^2}dx\] which gave \[-\frac{1}{1+e^r}+\frac{1}{2}\] by plugging in upper limit and lower limit into the anti derivative

OpenStudy (anonymous):

in other words just like you would for a "regular " integral, once you compute the anti derivative \[F(x)\] you find \[F(b)-F(a)\]

OpenStudy (anonymous):

okay so when did we double it?

OpenStudy (anonymous):

(btw I plugged in 0 and b and got the same things)

OpenStudy (anonymous):

at the very end. you get 1/2 as your integral, double it and get 1

OpenStudy (anonymous):

b, r, whatever your book uses and you like

OpenStudy (anonymous):

most actually use l

OpenStudy (anonymous):

okay so after I find the integral from 0 to b... why do I just have 1/2?

OpenStudy (anonymous):

the integral from 0 to b gave you \[-\frac{1}{1+e^b}+\frac{1}{2}\] right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now you take the limit as b goes to infinity, because that is the definition of the "improper" integral

OpenStudy (anonymous):

\[\int_0^{\infty} \frac{e^x}{(1+e^x)^2}dx=\lim_{b\rightarrow \infty}\int_0^b \frac{e^x}{(1+e^x)^2}dx=\lim_{b\rightarrow \infty}-\frac{1}{1+e^b}+\frac{1}{2}=0+\frac{1}{2}\]

OpenStudy (anonymous):